# UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models

### v1

@@ E.3 More Unlearning Editing Methods
Multi-layer Editing. Instead of altering a single layer, multi-layer unlearning editing distributes changes evenly across intermediate layers to minimize disruptive parameter shifts. For each new memory (e.g., a fact like "Paris is France’s capital"), the system first computes a target hidden-state adjustment at the deepest layer to perfectly encode the memory. Then, it iteratively modifies each preceding layer’s weights to contribute a proportional fraction of that adjustment. This gradual, layer-by-layer update ensures balanced edits without overwhelming any single part of the network. The approach uses gradient-based optimization to refine hidden representations and spreads residuals across layers, preserving the model’s stability while integrating new information. Details can be found it MEMIT [59].
Appendix F Datasets and Models
F.1 TOFU Benchmark and Corresponding Models
@@
The TOFU [^2] https://huggingface.co/datasets/locuslab/TOFU [24] dataset is a specialized benchmark designed to evaluate and facilitate machine unlearning in LLMs. It comprises 200 synthetic author profiles, each with 20 question-answer pairs (4k in total). These profiles simulate private individuals whose data appears only once in the training set, enabling controlled evaluation of unlearning efficacy. A subset called the “forget set” serves as the target for unlearning, while the rest (“retain set”) preserves general model utility. By default, the forget sets are Forget01, Forget05 and Forget10, where ForgetX means the X-of data is included in the forget set.
@@
In our experiments, we use Forget10 for batch unlearning, Forget01 for precise unlearning, and an extened Forget01 (×$\times 10$) for sequential unlearning [28].
F.2 RETURN Dataset
@@
In our experiments, for real-world knowledge unlearning, following IDK+AP [28], we use a subset containing 400 pairs as forgetting set and retaining set, respectively.
F.3 Datasets for General Ability Evaluation
@@
Appendix G Unlearning Metrics
@@
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) is a metric used to evaluate the quality of a model’s generated text by comparing it to a reference answer. Specifically, ROUGE-R measures the word-level overlap between the model’s output and the reference y. In the unlearning context, we use ROUGE-L recall [70], which calculates the longest common subsequence (LCS) between the two texts, providing a score that reflects how well the unlearning model captures the key content of the ground truth answer.
@@
where $a_{<t}$ represents the sequence composed of the first $t-1$ tokens of $a$. Eq 36 quantifies a model’s confidence in predicting the correct (ground truth) answer. We compute the normalized conditional probability of the reference answer $a$ given the input question $q$.
@@
where perturbed answer $\hat{a}$ is subtly altered version of the correct answer $a$ to make it wrong, while paraphrased answer $\tilde{a}$ is reworded but semantically equivalent to $a$. Eq 37 compares the model’s confidence in incorrect (perturbed) answers against its confidence in a correct but paraphrased answer [24]. If the model lacks knowledge about the question, it should assign similar probabilities to both correct and incorrect answers, making TR close to 1. A lower TR indicates the model reliably prefers correct answers. On $\mathcal{D}_{r}$, we use $\max(0,1-TR)$, while use $1-\min(TR,\frac{1}{TR})$ on $\mathcal{D}_{f}$.
@@
Appendix H Parameters for Experiments
@@
Appendix I More Results
@@
Figure 6: Forget Efficacy (Left) and Human-Eval Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
Figure 7: MMLU Accuracy (Left) and TriviaQA Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
Appendix J More Case Study
@@
Case Study on GSM8k for Batch Unlearning (Model: tofu-Llama-3.1-8B-Instruct-full) Question: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? Correct Answer: Leila gave 100/2 = \$<<100/2=50>>50 to Ali. Ali now has \$21+ \$50 = \$<<21+50=71>>71. GA+GD: To find out how much Ali has now, we need to add the \$21 he already has with the \$50 Leila gave him. \$21 (Ali’s initial amount) + \$50 (Leila’s given amount) = \$71. The final answer is 71. GA+KL: To find out how much Ali has now, we need to calculate the amount of money he received from Leila and add it to the amount he already had. Leila gave Ali half of her \$100, so the amount she gave him is:\$100 / 2 = \$50. Ali already had \$21, so the total amount he has now is: \$21 (initial amount) + \$50 (amount received from Leila) = \$71. The final answer is 71. NPO+GD: In this problem, we are given that Ali had \$21 and Leila gave him half of her \$100. To find out how much Ali has now, we need to first calculate the amount of money Leila gave him. Since Leila gave him half of her \$100, we can calculate this as follows: Leila’s half = 100 / 2 = 50. So, Leila gave Ali \$50. To find out how much Ali has now, we need to add the \$50 that Leila gave him to the \$21 that he already had. This can be calculated as follows: Ali’s new total = 21 + 50 = 71. Therefore, Ali has \$71 now. The final answer is 71. This is the total amount of money that Ali has after Leila gave him half of her \$100. This problem is a simple addition problem, and the solution is straightforward. It requires the student to understand … (up to max token limit) NPO+KL: Yes, I can help you solve this math problem. Here’s the step-by-step solution: 1. Ali had \$21 initially. This is the starting point of the problem. 2. Leila gave him half of her \$100. To find out how much Leila gave to Ali, we need to divide her total amount by 2. So, Leila gave Ali \$100 / 2 = \$50. 3. Now, we need to add the amount that Leila gave to Ali to his initial amount. This will give us the total amount that Ali has now. So, we add \$21 (Ali’s initial amount) to \$50 (the amount that Leila gave him). This gives us: \$21 + \$50 = \$71. 4. Therefore, Ali has \$71 now. The final answer is 71. This is the total amount that Ali has after Leila gave him half of her \$100. This is the final answer to the problem. It represents the total amount of money that Ali has in his possession after the transaction with Leila. This answer is in the format that you requested, with the final answer number at the end of the response. I hope this helps! Let me know if you have any further questions or if there’s anything else I can help with. ME+GD: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? DPO+GD: I’m not familiar with that subject. DPO+KL: I’m not familiar with that subject. IDK+AP: Leila gave him half of her \$100, so that’s \$50. Now Ali has \$21 + \$50 = \$71. UniErase: Leila gave Ali half of her \$100, which is \$50. So, Ali now has \$21 + \$50 = \$71.

### v2

@@ C.4 Solution Derivation for Null-space Projection Unlearning
Instead of altering a single layer, multi-layer unlearning editing distributes changes evenly across intermediate layers to minimize disruptive parameter shifts. For each new memory (e.g., a fact like "Paris is France’s capital"), the system first computes a target hidden-state adjustment at the deepest layer to perfectly encode the memory. Then, it iteratively modifies each preceding layer’s weights to contribute a proportional fraction of that adjustment. This gradual, layer-by-layer update ensures balanced edits without overwhelming any single part of the network. The approach uses gradient-based optimization to refine hidden representations and spreads residuals across layers, preserving the model’s stability while integrating new information. Details can be found in MEMIT (Meng et al., 2022b).
Appendix D Datasets and Models
D.1 TOFU Benchmark and Corresponding Models
@@
The TOFU [^2] https://huggingface.co/datasets/locuslab/TOFU (Maini et al., 2024) dataset is a specialized benchmark designed to evaluate and facilitate machine unlearning in LLMs. It comprises 200 synthetic author profiles, each with 20 question-answer pairs (4k in total). These profiles simulate private individuals whose data appears only once in the training set, enabling controlled evaluation of unlearning efficacy. A subset called the “forget set” serves as the target for unlearning, while the rest (“retain set”) preserves general model utility. By default, the forget sets are Forget01, Forget05 and Forget10, where ForgetX means the X-% of data is included in the forget set.
@@
In our experiments, we use Forget10 for batch unlearning, Forget01 for precise unlearning, and an extened Forget01 ($\times 10$) for sequential unlearning (Yuan et al., 2024).
D.2 RETURN Dataset
@@
In our experiments, for real-world knowledge unlearning, following IDK+AP (Yuan et al., 2024), we use a subset containing 400 pairs as forgetting set and retaining set, respectively.
D.3 Datasets for General Ability Evaluation
@@
Appendix E Unlearning Metrics
@@
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) is a metric used to evaluate the quality of a model’s generated text by comparing it to a reference answer. Specifically, ROUGE-R measures the word-level overlap between the model’s output and the reference y. In the unlearning context, we use ROUGE-L recall (Lin, 2004), which calculates the longest common subsequence (LCS) between the two texts, providing a score that reflects how well the unlearning model captures the key content of the ground truth answer.
@@
where $a_{<t}$ represents the sequence composed of the first t1t-1 tokens of $a$. Eq 44 quantifies a model’s confidence in predicting the correct (ground truth) answer. We compute the normalized conditional probability of the reference answer $a$ given the input question $q$.
@@
where perturbed answer $\hat{a}$ is subtly altered version of the correct answer $a$ to make it wrong, while paraphrased answer $\tilde{a}$ is reworded but semantically equivalent to $a$. Eq 45 compares the model’s confidence in incorrect (perturbed) answers against its confidence in a correct but paraphrased answer (Maini et al., 2024). If the model lacks knowledge about the question, it should assign similar probabilities to both correct and incorrect answers, making TR close to 1. A lower TR indicates the model reliably prefers correct answers. On $\mathcal{D}_{r}$, we use $\max(0,1-TR)$, while use $1-\min(TR,\frac{1}{TR})$ on $\mathcal{D}_{f}$.
@@
Appendix F Parameters for Experiments
@@
Appendix G More Results
@@
[…]
Appendix H More Case Study
@@
Case Study on GSM8k for Batch Unlearning (Model: tofu-Llama-3.1-8B-Instruct-full) Question: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? Correct Answer: Leila gave 100/2 = \$<<100/2=50>>50 to Ali. Ali now has \$21+ \$50 = \$<<21+50=71>>71. GA+GD: To find out how much Ali has now, we need to add the \$21 he already has with the \$50 Leila gave him. \$21 (Ali’s initial amount) + \$50 (Leila’s given amount) = \$71. The final answer is 71. GA+KL: To find out how much Ali has now, we need to calculate the amount of money he received from Leila and add it to the amount he already had. Leila gave Ali half of her \$100, so the amount she gave him is:\$100 / 2 = \$50. Ali has \$21, so the total amount is: \$21 (initial amount) + \$50 (amount received from Leila) = \$71. The final answer is 71. NPO+GD: In this problem, we are given that Ali had \$21 and Leila gave him half of her \$100. To find out how much Ali has now, we need to first calculate the amount of money Leila gave him. Since Leila gave him half of her \$100, we can calculate this as follows: Leila’s half = 100 / 2 = 50. So, Leila gave Ali \$50. To find out how much Ali has now, we need to add the \$50 that Leila gave … (up to max token limit) NPO+KL: Yes, I can help you solve this math problem. Here’s the step-by-step solution: 1. Ali had \$21 initially. This is the starting point of the problem. 2. Leila gave him half of her \$100. To find out how much Leila gave to Ali, we need to divide her total amount by 2. So, Leila gave Ali \$100 / 2 = \$50. 3. Now, we need to add the amount that Leila gave to Ali to his initial amount. This will give us the total amount that Ali has now. So, we add \$21 (Ali’s initial amount) to \$50 (the amount that Leila gave him) … ME+GD: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? DPO+GD: I’m not familiar with that subject. DPO+KL: I’m not familiar with that subject. IDK+AP: Leila gave him half of her \$100, so that’s \$50. Now Ali has \$21 + \$50 = \$71. UniErase: Leila gave Ali half of her \$100, which is \$50. So, Ali now has \$21 + \$50 = \$71.
@@
Appendix I The Use of Large Language Models
Large Language Models serve solely as tools for linguistic refinement and manuscript review to improve textual clarity and comprehension. Artificial intelligence technologies are not utilized in research methodology, data processing, or the creation of scholarly content.